GNU CPIO

2.12 12 September 2015

by Robert Carleton and Sergey Poznyakoff

This manual documents GNU cpio (version 2.12, 12 September 2015). Copyright \begin{matrix}c \\ ◯ \end{matrix} 1995, 2001-2002, 2004, 2010, 2014-2015 Free Software Foundation, Inc.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.2 or any later version published by the Free Software Foundation; with no Invariant Sections, with the Front-Cover texts being “A GNU Manual”, and with the Back-Cover Texts as in (a) below. A copy of the license is included in the section entitled “GNU Free Documentation License”.

(a) The FSF’s Back-Cover Text is: “You have freedom to copy and modify this GNU Manual, like GNU software. Copies published by the Free Software Foundation raise funds for GNU development.”

Published by the Free Software Foundation 51 Franklin Street, Fifth Floor, Boston, MA 02110-1301, USA

1 Introduction

GNU cpio copies files into or out of a cpio or tar archive, The archive can be another file on the disk, a magnetic tape, or a pipe.

GNU cpio supports the following archive formats: binary, old ASCII, new ASCII, crc, HPUX binary, HPUX old ASCII, old tar, and POSIX.1 tar. The tar format is provided for compatibility with the tar program. By default, cpio creates binary format archives, for compatibility with older cpio programs. When extracting from archives, cpio automatically recognizes which kind of archive it is reading and can read archives created on machines with a different byte-order.

2 Tutorial

GNU cpio performs three primary functions. Copying files to an archive, Extracting files from an archive, and passing files to another directory tree. An archive can be a file on disk, one or more floppy disks, or one or more tapes.

When creating an archive, cpio takes the list of files to be processed from the standard input, and then sends the archive to the standard output, or to the device defined by the ‘-F’ option. See Section 3.1 [Copy-out mode], page 2. Usually find or ls is used to provide this list to the standard input. In the following example you can see the possibilities for archiving the contents of a single directory.

      % ls | cpio -ov > directory.cpio

The ‘-o’ option creates the archive, and the ‘-v’ option prints the names of the files archived as they are added. Notice that the options can be put together after a single ‘-’ or can be placed separately on the command line. The ‘>’ redirects the cpio output to the

file ‘directory.cpio’.

If you wanted to archive an entire directory tree, the find command can provide the file list to cpio:

      % find . -print -depth | cpio -ov > tree.cpio

This will take all the files in the current directory, the directories below and place them in the archive tree.cpio. Again the ‘-o’ creates an archive, and the ‘-v’ option shows you the name of the files as they are archived. See Section 3.1 [Copy-out mode], page 2. Using the ‘.’ in the find statement will give you more flexibility when doing restores, as it will save file names with a relative path vice a hard wired, absolute path. The ‘-depth’ option forces ‘find’ to print of the entries in a directory before printing the directory itself. This limits the effects of restrictive directory permissions by printing the directory entries in a directory before the directory name itself.

Extracting an archive requires a bit more thought because cpio will not create directories by default. Another characteristic, is it will not overwrite existing files unless you tell it to.

      % cpio -iv < directory.cpio

This will retrieve the files archived in the file directory.cpio and place them in the present directory. The ‘-i’ option extracts the archive and the ‘-v’ shows the file names as they are extracted. If you are dealing with an archived directory tree, you need to use the ‘-d’ option to create directories as necessary, something like:

      % cpio -idv < tree.cpio

This will take the contents of the archive tree.cpio and extract it to the current directory. If you try to extract the files on top of files of the same name that already exist (and have the same or later modification time) cpio will not extract the file unless told to do so by the -u option. See Section 3.2 [Copy-in mode], page 4.

In copy-pass mode, cpio copies files from one directory tree to another, combining the copy-out and copy-in steps without actually using an archive. It reads the list of files to copy from the standard input; the directory into which it will copy them is given as a non-option argument. See Section 3.3 [Copy-pass mode], page 6.

      % find . -depth -print0 | cpio --null -pvd new-dir

The example shows copying the files of the present directory, and sub-directories to a new directory called new-dir. Some new options are the ‘-print0’ available with GNU find, combined with the ‘--null’ option of cpio. These two options act together to send file names between find and cpio, even if special characters are embedded in the file names. Another is ‘-p’, which tells cpio to pass the files it finds to the directory ‘new-dir’.

3 Invoking cpio

3.1 Copy-out mode

In copy-out mode, cpio copies files into an archive. It reads a list of filenames, one per line, on the standard input, and writes the archive onto the standard output. A typical way to generate the list of filenames is with the find command; you should give find the ‘-depth’ option to minimize problems with permissions on directories that are unreadable.

Copy-out mode is requested by the ‘-o’ (‘--create’) command line option, e.g.:

      % find | cpio -o > directory.cpio

The following options can be used in copy-out mode:

‘-0’

‘--null’ Filenames in the list are delimited by ASCII null characters instead of newlines.

‘-A’
‘--append’

Append to an existing archive.

‘-a’
‘--reset-access-time’

Reset the access times of files after reading them.

‘--absolute-filenames’

Do not strip file system prefix components from the file names.

‘--no-absolute-filenames’

Strip file system prefix components from the file names before storing them to the archive.

‘--block-size=block-size’

Sets the I/O block size to block-size * 512 bytes.

‘-B’ Set the I/O block size to 5120 bytes.

‘-c’ Use the old portable (ASCII) archive format.

‘-C number’
‘--io-size=number’

Set the I/O block size to the given number of bytes.

‘-D dir’
‘--directory=dir’

Change to directory dir

‘--force-local’

Treat the archive file as local, even if its name contains colons.

‘-F [[user@]host:]archive-file’
‘-O [[user@]host:]archive-file’
‘--file=[[user@]host:]archive-file’

Use the supplied archive-file instead of standard input. Optional user and host specify the user and host names in case of a remote archive.

‘-H format’
‘--format=format’

Use given archive format. See [format], page 10, for a list of available formats.

‘-L’
‘--dereference’

Dereference symbolic links (copy the files that they point to instead of copying the links).

‘-M string’
‘--message=string’

Print string when the end of a volume of the backup media is reached.

‘--quiet’ Do not print the number of blocks copied.

‘--rsh-command=command’

Use command instead of rsh to access remote archives.

‘-R’
‘--owner=[user][:.][group]’

Set the ownership of all files created to the specified user and/or group. See [owner], page 12.

‘-v’
‘--verbose’

Verbosely list the files processed.

‘-V’

‘--dot’ Print a ‘.’ for each file processed.

‘-W’
‘--warning=flag’

Control warning display. Argument is one of ‘none’, ‘truncate’, ‘no-truncate’ or ‘all’. See [warning], page 14, for a detailed discussion of these.

3.2 Copy-in mode

In copy-in mode, cpio copies files from an archive into the filesystem or lists the archive contents. It reads the archive from the standard input. Any non-option command line arguments are shell globbing patterns; only files in the archive whose names match one or more of those patterns are copied from the archive. Unlike in the shell, an initial ‘.’ in a filename does match a wildcard at the start of a pattern, and a ‘/’ in a filename can match wildcards. If no patterns are given, all files are extracted.

The copy-in mode is requested by the ‘-i’ (‘--extract’) command line option.

The following options can be used in copy-in mode:

‘--absolute-filenames’

Do not strip file system prefix components from the file names.

‘--no-absolute-filenames’

Create all files relative to the current directory.

‘--block-size=block-size’

Sets the I/O block size to block-size * 512 bytes.

‘-b’

‘--swap’ Swap both halfwords of words and bytes of halfwords in the data. Equivalent to ‘-sS’.

‘-B’ Set the I/O block size to 5120 bytes.

‘-c’ Use the old portable (ASCII) archive format.

‘-C number’
‘--io-size=number’

Set the I/O block size to the given number of bytes.

‘-D dir’
‘--directory=dir’

Change to directory dir

‘-d’
‘--make-directories’

Create leading directories where needed.

‘-E file’
‘--pattern-file=file’

Read additional patterns specifying filenames to extract or list from file.

‘-f’
‘--nonmatching’

Only copy files that do not match any of the given patterns.

‘--force-local’

Treat the archive file as local, even if its name contains colons.

‘-F [[user@]host:]archive-file’
‘-I [[user@]host:]archive-file’
‘--file=[[user@]host:]archive-file’

Use the supplied archive-file instead of standard input. Optional user and host specify the user and host names in case of a remote archive.

‘-H format’
‘--format=format’

Use given archive format. See [format], page 10, for a list of available formats.

‘-m’
‘--preserve-modification-time’

Retain previous file modification times when creating files.

‘-M string’
‘--message=string’

Print string when the end of a volume of the backup media is reached.

‘--no-preserve-owner’

Do not change the ownership of the files.

‘-n’
‘--numeric-uid-gid’

In the verbose table of contents listing, show numeric UID and GID values.

‘--only-verify-crc’

When reading a CRC format archive, only verify the CRC’s of each file in the archive, don’t actually extract the files

‘--quiet’ Do not print the number of blocks copied.

‘--rsh-command=command’

Use command instead of rsh to access remote archives.

‘-r’
‘--rename’

Interactively rename files

‘--sparse’

Write files with large blocks of zeros as sparse files.

‘-s’
‘--swap-bytes’

Swap the bytes of each halfword in the files

‘-S’
‘--swap-halfwords’

Swap the halfwords of each word (4 bytes) in the files

‘--to-stdout’

Extract files to standard output.

‘-u’
‘--unconditional’

Replace all files unconditionally.

‘-v’
‘--verbose’

Verbosely list the files processed.

‘-V’

‘--dot’ Print a ‘.’ for each file processed.

‘-W’
‘--warning=flag’

Control warning display. Argument is one of ‘none’, ‘truncate’, ‘no-truncate’ or ‘all’. See [warning], page 14, for a detailed discussion of these.

3.3 Copy-pass mode

In copy-pass mode, cpio copies files from one directory tree to another, combining the copyout and copy-in steps without actually using an archive. It reads the list of files to copy from the standard input; the directory into which it will copy them is given as a non-option argument.

This mode is requested by the ‘-p’ (‘--pass-through’) command line option.

The following options are valid in copy-out mode:

‘-0’

‘--null’ Filenames in the list are delimited by ASCII null characters instead of newlines.

‘-a’
‘--reset-access-time’

Reset the access times of files after reading them.

‘-b’

‘--swap’ Swap both halfwords of words and bytes of halfwords in the data. Equivalent to ‘-sS’.

‘--block-size=block-size’

Sets the I/O block size to block-size * 512 bytes.

‘-B’ Set the I/O block size to 5120 bytes.

‘-c’ Use the old portable (ASCII) archive format.

‘-C number’
‘--io-size=number’

Set the I/O block size to the given number of bytes.

‘-d’
‘--make-directories’

Create leading directories where needed.

‘--device-independent’
‘--reproducible’

Create reproducible archives. This is equivalent to ‘--ignore-devno

          --renumber-inodes’.
‘-D dir’
‘--directory=dir’

Change to directory dir

‘-E file’
‘--pattern-file=file’

Read additional patterns specifying filenames to extract or list from file.

‘-f’
‘--nonmatching’

Only copy files that do not match any of the given patterns.

‘-F [[user@]host:]archive-file’
‘-O [[user@]host:]archive-file’
‘--file=[[user@]host:]archive-file’

Use the supplied archive-file instead of standard input. Optional user and host specify the user and host names in case of a remote archive.

‘--force-local’

Treat the archive file as local, even if its name contains colons.

‘-H format’
‘--format=format’

Use given archive format. See [format], page 10, for a list of available formats.

‘--ignore-devno’

Store 0 in the device number field of each archive member, instead of the actual device number.

‘-l’

‘--link’ Link files instead of copying them, when possible.

‘-L’
‘--dereference’

Dereference symbolic links (copy the files that they point to instead of copying the links).

‘-m’
‘--preserve-modification-time’

Retain previous file modification times when creating files.

‘-M string’
‘--message=string’

Print string when the end of a volume of the backup media is reached.

‘-n’
‘--numeric-uid-gid’

In the verbose table of contents listing, show numeric UID and GID values.

‘--no-preserve-owner’

Do not change the ownership of the files.

‘--only-verify-crc’

When reading a CRC format archive, only verify the CRC’s of each file in the archive, don’t actually extract the files

‘--quiet’ Do not print the number of blocks copied.

‘--rsh-command=command’

Use command instead of rsh to access remote archives.

‘-r’
‘--rename’

Interactively rename files

‘--renumber-inodes’

Renumber inodes when storing them in the archive.

‘-R’
‘--owner=[user][:.][group]’

Set the ownership of all files created to the specified user and/or group. See [owner], page 12.

‘-s’
‘--swap-bytes’

Swap the bytes of each halfword in the files

‘--sparse’

Write files with large blocks of zeros as sparse files.

‘-S’
‘--swap-halfwords’

Swap the halfwords of each word (4 bytes) in the files

‘--to-stdout’

Extract files to standard output.

‘-u’
‘--unconditional’

Replace all files unconditionally.

‘-v’
‘--verbose’

Verbosely list the files processed.

‘-V’

‘--dot’ Print a ‘.’ for each file processed.

‘-W’

‘--warning=flag’

Control warning display. Argument is one of ‘none’, ‘truncate’, ‘no-truncate’ or ‘all’. See [warning], page 14, for a detailed discussion of these.

3.4 Options

This section summarizes all available command line options. References in square brackets after each option indicate cpio modes in which this option is valid.

-0

--null [[copy-in], page 4,[copy-out], page 2,[copy-pass], page 6] Read a list of filenames terminated by a null character, instead of a newline, so that files whose names contain newlines can be archived. GNU find is one way to produce a list of null-terminated filenames. This option may be used in copy-out and copy-pass modes.

-a
--reset-access-time

[[copy-out], page 2,[copy-pass], page 6] Reset the access times of files after reading them, so that it does not look like they have just been read.

-A

--append [[copy-out], page 2] Append to an existing archive. Only works in copy-out mode. The archive must be a disk file specified with the ‘-O’ or ‘-F’ (‘--file’) option.

-b

--swap [[copy-in], page 4] Swap both halfwords of words and bytes of halfwords in the data. Equivalent to ‘-sS’. This option may be used in copy-in mode. Use this option to convert 32-bit integers between big-endian and little-endian machines.

-B [[copy-in], page 4,[copy-out], page 2,[copy-pass], page 6] Set the I/O block size to 5120 bytes. Initially the block size is 512 bytes.

--block-size=block-size

[[copy-in], page 4,[copy-out], page 2,[copy-pass], page 6] Set the I/O block size to block-size * 512 bytes.

-c [[copy-in], page 4,[copy-out], page 2,[copy-pass], page 6] Use the old portable (ASCII) archive format.

-C io-size
--io-size=io-size

[[copy-in], page 4,[copy-out], page 2,[copy-pass], page 6] Set the I/O block size to io-size bytes.

-d
--make-directories

[[copy-in], page 4,[copy-pass], page 6] Create leading directories where needed.

-D dir
--directory=dir

[[copy-in], page 4,[copy-out], page 2,[copy-pass], page 6] Change to the directory dir before starting the operation. This can be used, for example, to extract an archive contents in a different directory:

     $ cpio -i -D /usr/local < archive

or to copy-pass files from one directory to another:

     $ cpio -D /usr/bin -p /usr/local/bin < filelist

The ‘-D’ option does not affect file names supplied as arguments to another command line options, such as ‘-F’ or ‘-E’. For example, the following invocation:

     cpio -D /tmp/foo -d -i -F arc

instructs cpio to open the archive file ‘arc’ in the current working directory, then change to the directory ‘/tmp/foo’ and extract files to that directory. If ‘/tmp/foo’ does not exist, it will be created first (the ‘-d’ option) and then changed to.

-E file
--pattern-file=file

[[copy-in], page 4] Read additional patterns specifying filenames to extract or list from file. The lines of file are treated as if they had been non-option arguments to cpio. This option is used in copy-in mode,

-f
--nonmatching

[[copy-in], page 4] Only copy files that do not match any of the given patterns.

-F archive
--file=archive

[[copy-in], page 4,[copy-out], page 2] Archive filename to use instead of standard input or output. To use a tape drive on another machine as the archive, use a filename that starts with ‘hostname:’, where hostname is the name or IP address of the machine. The hostname can be preceded by a username and an ‘@’ to access the remote tape drive as that user, if you have permission to do so (typically an entry in that user’s ‘~/.rhosts’ file).

--force-local

[[copy-in], page 4,[copy-out], page 2] With ‘-F’, ‘-I’, or ‘-O’, take the archive file name to be a local file even if it contains a colon, which would ordinarily indicate a remote host name.

-H format
--format=format

[[copy-in], page 4,[copy-out], page 2,[copy-pass], page 6] Use archive format format. The valid formats are listed below with file size limits for individual files in parentheses; the same names are also recognized in all-caps. The default in copy-in mode is to automatically detect the archive format, and in copy-out mode is ‘bin’.

‘bin’ The obsolete binary format. (2147483647 bytes)

‘odc’ The old (POSIX.1) portable format. (8589934591 bytes)

‘newc’ The new (SVR4) portable format, which supports file systems having more than 65536 i-nodes. (4294967295 bytes)

‘crc’ The new (SVR4) portable format with a checksum added.

‘tar’ The old tar format. (8589934591 bytes)

‘ustar’ The POSIX.1 tar format. Also recognizes GNU tar archives, which are similar but not identical. (8589934591 bytes)

‘hpbin’ The obsolete binary format used by HPUX’s cpio (which stores device files differently).

‘hpodc’ The portable format used by HPUX’s cpio (which stores device files differently).

-i
--extract

Run in copy-in mode. See Section 3.2 [Copy-in mode], page 4.

-I archive

[[copy-in], page 4] Archive filename to use instead of standard input. To use a tape drive on another machine as the archive, use a filename that starts with ‘hostname:’, where hostname is the name or IP address of the remote host. The hostname can be preceded by a username and an ‘@’ to access the remote tape drive as that user, if you have permission to do so (typically an entry in that user’s

‘~/.rhosts’ file).
-l

--link [[copy-pass], page 6] Link files instead of copying them, when possible.

-L
--dereference

[[copy-in], page 4,[copy-pass], page 6] Copy the file that a symbolic link points to, rather than the symbolic link itself.

-m
--preserve-modification-time

[[copy-in], page 4,[copy-pass], page 6] Retain previous file modification times when creating files.

-M message
--message=message

[[copy-in], page 4,[copy-out], page 2] Print message when the end of a volume of the backup media (such as a tape or a floppy disk) is reached, to prompt the user to insert a new volume. If message contains the string ‘%d’, it is replaced by the current volume number (starting at 1).

-n
--numeric-uid-gid

[[copy-in], page 4] Show numeric UID and GID instead of translating them into names when using the ‘--verbose’ option.

--no-absolute-filenames

[[copy-in], page 4,[copy-out], page 2] Create all files relative to the current directory in copy-in mode, even if they have an absolute file name in the archive.

--no-preserve-owner

[[copy-in], page 4,[copy-pass], page 6] Do not change the ownership of the files; leave them owned by the user extracting them. This is the default for non-root users, so that users on System V don’t inadvertantly give away files. This option can be used in copy-in mode and copy-pass mode

-o

--create Run in copy-out mode. See Section 3.1 [Copy-out mode], page 2.

-O archive

[[copy-out], page 2] Archive filename to use instead of standard output. To use a tape drive on another machine as the archive, use a filename that starts with ‘hostname:’, where hostname is the name or IP address of the machine. The hostname can be preceded by a username and an ‘@’ to access the remote tape drive as that user, if you have permission to do so (typically an entry in that user’s

‘~/.rhosts’ file).
--only-verify-crc

[[copy-in], page 4] Verify the CRC’s of each file in the archive, when reading a CRC format archive. Don’t actually extract the files.

-p
--pass-through

Run in copy-pass mode. See Section 3.3 [Copy-pass mode], page 6.

--quiet [[copy-in], page 4,[copy-out], page 2,[copy-pass], page 6] Do not print the number of blocks copied.

-r

--rename [[copy-in], page 4] Interactively rename files.

-R owner
--owner owner

[[copy-in], page 4,[copy-out], page 2,[copy-pass], page 6] In copy-in and copy-pass mode, set the ownership of all files created to the specified owner (this operation is allowed only for the super-user). In copy-out mode, store the supplied owner information in the archive. The argument can be either the user name or the user name and group name, separated by a dot or a colon, or the group name, preceeded by a dot or a colon, as shown in the examples below:

      cpio --owner smith
      cpio --owner smith:
      cpio --owner smith:users
      cpio --owner :users

The argument parts are first looked up in the system user and group databases, correspondingly. If any of them is not found there, it is treated as numeric UID or GID, provided that it consists of decimal digits only. To avoid the lookup and ensure that arguments are treated as numeric values, prefix them with a plus sign, e.g.:

      cpio --owner +0
      cpio --owner +0:
      cpio --owner +0:+0
      cpio --owner :+0

If the group is omitted but the ‘:’ or ‘.’ separator is given, as in the second example. the given user’s login group will be used.

--rsh-command=command

[[copy-in], page 4,[copy-out], page 2,[copy-pass], page 6] Notifies cpio that is should use command to communicate with remote devices.

-s
--swap-bytes

[[copy-in], page 4] Swap the bytes of each halfword (pair of bytes) in the files. This option can be used in copy-in mode.

-S
--swap-halfwords

[[copy-in], page 4] Swap the halfwords of each word (4 bytes) in the files. This option may be used in copy-in mode.

--sparse [[copy-in], page 4,[copy-pass], page 6] Write files with large blocks of zeros as sparse files. This option is used in copy-in and copy-pass modes.

-t

--list [[copy-in], page 4] Print a table of contents of the input. Can be used alone, as a mode designator, in which case ‘-i’ is implied.

--to-stdout

[[copy-in], page 4]

Extract files to standard output. This option may be used in copy-in mode.

-u
--unconditional

[[copy-in], page 4,[copy-pass], page 6]

Replace all files, without asking whether to replace existing newer files with older files.

-v
--verbose

[[copy-in], page 4,[copy-out], page 2,[copy-pass], page 6]

List the files processed, or with ‘-t’, give an ‘ls -l’ style table of contents listing. In a verbose table of contents of a ustar archive, user and group names in the archive that do not exist on the local system are replaced by the names that correspond locally to the numeric UID and GID stored in the archive.

-V

--dot [[copy-in], page 4,[copy-out], page 2,[copy-pass], page 6] Print a ‘.’ for each file processed.

--version

Print the cpio program version number and exit.

-W
--warning=flag

[[copy-in], page 4,[copy-out], page 2,[copy-pass], page 6]

Control warning display. The argument is one of the following:

none Disable all warnings.

all Enable all warnings.

truncate Warn about truncation of file header fields.

no-truncate Disable truncation warnings.

4 Magnetic Media

Archives are usually written on removable media–tape cartridges, mag tapes, or floppy disks.

The amount of data a tape or disk holds depends not only on its size, but also on how it is formatted. A 2400 foot long reel of mag tape holds 40 megabytes of data when formated at 1600 bits per inch. The physically smaller EXABYTE tape cartridge holds 2.3 gigabytes.

Magnetic media are re-usable–once the archive on a tape is no longer needed, the archive can be erased and the tape or disk used over. Media quality does deteriorate with use, however. Most tapes or disks should be disgarded when they begin to produce data errors.

Magnetic media are written and erased using magnetic fields, and should be protected from such fields to avoid damage to stored data. Sticking a floppy disk to a filing cabinet using a magnet is probably not a good idea.

Concept Index 15

5 Reporting bugs or suggestions

It is possible you will encounter a bug in cpio. If this happens, we would like to hear about it. As the purpose of bug reporting is to improve software, please be sure to include maximum information when reporting a bug. The information needed is:

• Version of the package you are using.

• Compilation options used when configuring the package.

• Conditions under which the bug appears.

Send your report to <bug-cpio@gnu.org>. Allow us a couple of days to answer.

Concept Index

A

archive creation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 archive extraction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

C

command line options . . . . . . . . . . . . . . . . . . . . . . . . . 2 copy files between filesystems . . . . . . . . . . . . . . . . . . 6 copy-in. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 copy-out . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 copy-pass . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 copying directory structures . . . . . . . . . . . . . . . . . . . 1 create archive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 creating a cpio archive . . . . . . . . . . . . . . . . . . . . . . . . 1

E

extract files from an archive . . . . . . . . . . . . . . . . . . . 4 extracting a cpio archive . . . . . . . . . . . . . . . . . . . . . . 1

I

invoking cpio . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

M

magnetic media. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

P

passing directory structures. . . . . . . . . . . . . . . . . . . . 1

Table of Contents

1 Introduction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 2 Tutorial . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 3 Invoking cpio . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

3.1 Copy-out mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 3.2 Copy-in mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 3.3 Copy-pass mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 3.4 Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

4 Magnetic Media . . . . . . . . . . . . . . . . . . . . . . . . . . 14 5 Reporting bugs or suggestions . . . . . . . . . . . . . 15 Concept Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15